# Fast SDE-based Monte Carlo dose calculation for proton therapy validated against Geant4

### v1

@@ Abstract
Objective: To validate a newly proposed stochastic differential equation (SDE)-based model for proton beam energy deposition by comparing its predictions with those from Geant4 in simplified phantom scenarios.
Approach: Building on previous work in SDE, where energy deposition from a proton beam was modelled using an SDE framework, we implemented the model with standard approximations to interaction cross sections and mean excitation energies, which makes simulations easily adaptable to new materials and configurations. The model was benchmarked against Geant4 in homogeneous and heterogeneous phantoms.
Main results: The SDE-based dose distributions agreed well with Geant4, showing range differences within 0.4 mm and 3D gamma pass rates exceeding 98% under 3%/2 mm criteria with a 1% dose threshold. The model achieved a computational speed-up of approximately fivefold relative to Geant4, consistent across different Geant4 physics lists.
Significance: These results demonstrate that the SDE approach can reproduce accuracy comparable to high-fidelity Monte Carlo for proton therapy at a fraction of the computational cost, highlighting its potential for accelerating dose calculations and treatment planning.
@@ 1 Introduction
Accurate dose calculations are critical for proton therapy treatment planning to ensure precise treatment delivery while minimising uncertainties in patient outcomes. Thus, high-fidelity models are needed to capture the inherently stochastic nature of proton interactions in tissue, including energy loss, scattering, and nuclear interactions. Currently, pencil beam algorithms (hong1996pencil; schaffner1999dose) are the standard dose calculation technique in the clinical setting because of their computational efficiency, which allows for fast treatment planning. However, they achieve speed at the expense of accuracy. Pencil beam algorithms perform reliably in homogeneous media and provide good agreement for depth–dose profiles, but struggle to accurately model lateral spread, tissue heterogeneities, and complex geometries. Due to this, Monte Carlo (MC) models such as Geant4 (agostinelli2003geant4), TOPAS (perl2012topas), GATE (grevillot2020gate) and FLUKA (bohlen2014fluka), among others, are an alternative to analytical methods to perform the most accurate dose calculations, which can reduce the error margins by several millimetres (paganetti2012range). This is achieved by tracking each proton step by step and sampling each possible interaction according to its probability, which means this method tends to be much slower, compromising computational efficiency and therefore its suitability for routine treatment planning.
GPU-based implementations such as MOQUI (lee2022moqui) enable highly efficient dose scoring, while other approaches avoid simulating secondary particles to fully exploit GPU computing power (shan2022virtual). On the CPU side, optimised codes such as MCsquare (deng2020mcsquare) leverage multi-core architectures to achieve substantial reductions in computation time. These developments are bringing MC closer to routine use in treatment planning and plan verification. However, achieving near real-time performance for adaptive therapy or robust optimisation remains challenging, particularly for complex patient geometries and medical images with a large number of voxels.
@@
To address this gap, new mathematical approaches are needed for modelling proton beam treatment delivery, aiming to lie between high-fidelity Monte Carlo accuracy and pencil beam algorithm efficiency. Tristan2 proposed an approach based on optimisation by computing the fluence of the associated Boltzmann transport equation numerically. In SDE, a complementary approach was introduced to solve the Boltzmann transport equation in backward form using a jump stochastic differential equation (SDE). The latter retains the probabilistic nature of Monte Carlo simulations while replacing frequent small-angle Coulomb scatterings with a physics-informed diffusion term. This reduces computational cost without sacrificing physical fidelity due to the central limit theorem. Moreover, KPP showed that the two methods are consistent in the sense that they are alternative but compatible mathematical descriptions of the same underlying dynamics, expressed in terms of macroscopic fluence on the one hand and microscopic particles on the other. The reduced runtime of the proposed SDE approach compared to full Monte Carlo simulations makes it an attractive candidate for tasks such as plan verification, robustness evaluation, and adaptive planning, where high accuracy is required but computational cost remains a limiting factor.
@@
  • To show SDE-based methods provide accurate dose predictions at low computational cost in clinically relevant settings when compared to current Monte Carlo codes.
@@
2 Microscopic proton physics
The fundamental basis of proton beam therapy is the principle of high energy protons decelerating through their interactions with subatomic particles in the surrounding tissue. Deceleration is synonymous with energy deposition, and this energy transfer is a key component of attacking cancerous tissues.
As the energy of a proton decreases, its rate of interaction with subatomic particles increases, leading to a greater rate of energy deposition. This means that, to a certain extent, the location at which the majority of the energy from a proton beam is deposited is controllable, causing minimal exposure to surrounding healthy tissues. Indeed, the energy deposition can be arranged in a way that there is no exit dose from a proton beam directed towards tumour tissues. These particular phenomena distinguish proton beam therapy from photon therapy.
In SDE, the basic principles of atomic physics were used to build a mathematical model for the dynamics of protons in a proton beam travelling through matter in three-dimensional space. The principal mathematical tool was a so-called stochastic differential equation (SDE) with jumps. The SDE method lends itself well to new Monte Carlo approaches which favourably balance accuracy and computational cost.
We begin with a brief reminder of the physics of proton beams and how this plays into the modelling choices that lead to the SDE formulation, which is introduced in the subsequent section. For a more detailed read on the background physics of proton beams, the reader is referred to NZ; Harald and Gott, among many possible sources.
2.1 Proton interactions
@@
In order to describe the evolution of protons, we need to introduce the space-direction-energy phase space in which they live. To that end, let $D\subset\mathbb{R}^{3}$ denote a closed and bounded convex spatial domain, $\mathbb{S}_{2}$ be the unit sphere in $\mathbb{R}^{3}$ and $\mathcal{E}=[\mathtt{e}_{\mathtt{min}},\mathtt{e}_{\mathtt{max}}]\subset[0,\infty)$ be the range of energies a proton can take. We thus define the energy-position-direction phase space as $\mathcal{C}=\mathcal{E}\times D\times\mathbb{S}_{2}.$
@@
Consider now a proton travelling through matter with an instantaneous configuration $x=({\mathrm{\epsilon}},r,\omega)\in\mathcal{C}$. In order to describe the dynamics of an individual proton, we consider the notion of transport and three broad types of atomic interactions with surrounding matter (See Figure 1). We refer the reader to NZ for further details. These mechanisms are:
@@
Transport. In the space between atoms, the proton moves in straight lines in the direction $\omega$.
@@
Inelastic Coulomb interaction. We assume that protons lose energy continuously as a consequence of a large number of collisions with orbital electrons. For an instantaneous configuration $x=({\mathrm{\epsilon}},r,\omega)$, there are two contributing factors to this continuous energy loss. Deterministic loss which occurs at a rate (per unit track length) given by what is commonly referred to as the stopping cross section, $\varsigma_{1}(x)$. We model variability in the number of orbital electron collisions per unit track length by an additional stochastic loss term, governed by a Brownian motion whose volatility is given by what is commonly referred to as the energy-straggling cross section $\varsigma_{2}(x)$. We assume that these collisions have no effect on the direction of motion of the relatively massive proton, or equivalently, that any effect is subsumed into the elastic scatter term described below.
@@
Figure 1: The three main interactions of a proton with matter. An elastic scattering (top) with the nucleus, a proton-nucleus collision which may be elastic or inelastic (centre), and an inelastic Coulomb interaction with atomic electrons (bottom).
@@
Elastic scatter. This class of events corresponds to a change in direction of the proton due to an interaction with a nearby nucleus whilst conserving energy between the proton and the nucleus. The three types of elastic scattering events we consider are as follows:
@@
Elastic Coulomb scatter: The proton passes sufficiently close to an atomic nucleus that it feels a repulsive positive charge, resulting in a change in its direction of motion. We distinguish between two cases: large and small scatters. For incoming direction $\omega$ and outgoing direction $\omega^{\prime}$, let $\theta\in[0,\pi]$ denote the polar scattering angle. Then, we fix a small cut-off, say $\delta>0$, such that the case $\theta>\delta$ is termed a large (elastic Coulomb) scatter and the case $\theta\leq\delta$ is called a small scatter. We use a cross section $\sigma_{\rm e}(x)\pi_{{\rm e}}(x;{\mathrm{d}}\omega^{\prime})$ to mathematically describe large scatter events, where $\sigma_{\rm e}(x)$ denotes the rate at which a proton with configuration $x=(\epsilon,r,\omega)\in\mathcal{C}$ undergoes a large elastic Coulomb scatter, and the probability density $\pi_{{\rm e}}(x;{\mathrm{d}}\omega^{\prime})$ describes the distribution of the resulting outgoing direction $\omega^{\prime}$. Small scatter events are subsumed into an aggregated term, described below. In the (medical) physics literature, it is not uncommon to think of $\sigma_{\rm e}(x)$ as a rate per unit track length, which captures the amount of atomic interactions a proton has with the medium through which it travels.
@@
Elastic proton-nucleus collision and scatter: In rare cases, the incoming proton is absorbed into the nucleus, which then becomes excited, resulting in a proton being released with the same energy as the incoming proton. Since protons are indistinguishable, we model them identically to the elastic Coulomb scatters described above, so that $\sigma_{e}$ and $\pi_{e}$ are the cross section and probability density of both event types combined.
@@
Coulomb-nuclear interference: It is insufficient to treat both the effects of elastic Coulomb scattering and elastic proton-nucleus collisions independently (Nekrasov_2024). The combination of these effects results in an additional perturbation in the outgoing angle due to so-called Coulomb-nuclear interference. We take the same cut-off $\delta$ to distinguish between small and large scatters, where large scatters are subsumed into the elastic Coulomb scatter and small scatters subsumed into the aggregation of small scatter, described below. Mathematically, $\sigma_{e}$ and $\pi_{e}$ incorporate Coulomb-nuclear interference as well.
@@
Aggregated small scatter. We consider the net effect of all small scatter events to result in a diffusive component in the direction of motion $\omega$ of protons (Vassiliev, Section 3.9.6). More precisely, the direction of transport $\omega$ undergoes a state-dependent Brownian motion on $\mathbb{S}_{2}$ with volatility $m(x)\geq 0$ where $x$ is the instantaneous configuration of the proton. As with large scatter events, we assume there is no energy loss associated with these small scattering events.
@@
Inelastic proton-nucleus collision and scatter. When a proton interacts with a nucleus, it can undergo an inelastic collision, forming an excited nuclear state. This excited nucleus may subsequently emit a spray of secondary particles, transferring some of its energy to each particle. For the SDE model considered in this article, we assume that each collision will emit exactly one secondary proton which carries the majority of the outgoing energy and outgoing direction $\omega^{\prime}$. The energy deficit between the incoming and outgoing protons is accounted for by the recoil of the nucleus, other secondary particle emission which our SDE neglects, and the binding energy of the interaction.
As will be made evident in Section 6, even with this secondary particle simplification, the accuracy of our model performs well against current Monte-Carlo codes without invoking the computational overhead incurred by these additional secondary tracks. Moreover, we emphasise that the SDE framework allows for multiple secondary particles, thus facilitating a hierarchical collection of models that trade computational cost for further accuracy. The result of this extension is a branching SDE, where SDEs of the form (1) are fitted to each particle type. However, for the scope of this article, we focus on how to fit a single type SDE for protons and leave this extension for future research.
Nuclear inelastic scatters occur according to the cross section $\sigma_{{\rm ne}}(x)$, seen as a rate per unit track length, in which case the configuration, $x$, of the incoming proton is transferred to an outgoing configuration $x^{\prime}=({\mathrm{\epsilon}}(1-u),r,\omega^{\prime})$ with probability density $\pi_{{\rm ne}}(x;{\mathrm{d}}\omega^{\prime},{\mathrm{d}}u)$, where $u\in(0,1]$.
2.2 Sequential proton tracks
In SDE, the notion of a sequential proton track was introduced to describe the evolution in configuration-space of an initial proton concatenated with subsequent protons that continue its trajectory after proton-nuclear interactions, ultimately to absorption or a de-energised state. The terminus point of the sequential proton track is called its range. Each sequential proton track is a random trajectory and we can think of a proton beam as being made up of many superimposed sequential proton tracks. Without loss of generality, it is assumed that each sequential proton track behaves independently and therefore, up to a scaling factor, by the Law of Large Numbers, the proton beam is nothing more than the average effect of an individual sequential proton track.
2.3 Bragg peak
@@
The Bragg peak is a graphical representation of the energy deposition per unit length along the axial direction of a proton beam. A typical Bragg peak curve is proportional to the average energy deposition per unit length along sequential proton tracks. Its characteristic shape, illustrated in Figure 2, shows a gradual increase in energy deposition followed by a sharp rise near the end of the proton range, and then a rapid fall-off. One of the defining features of the Bragg peak is its range, defined as the depth at which the dose falls to a predefined fraction of its maximum. Quantities such as R90 and R50, defined as the depths at which the dose falls to 90% and 50% of the maximum, respectively, are commonly used in the literature.
Figure 2: A typical Bragg Peak curve in a homogeneous medium. Depth is normalised to the proton range R90, and dose is normalised to its maximum value.
3 Sequential proton track SDE

### v2

# SDE-based Monte Carlo dose calculation for proton therapy validated against Geant4

@@ Abstract
Objective: To systematically assess the accuracy and computational performance of a newly proposed stochastic differential equation (SDE)-based model for proton beam dose calculation by benchmarking it against Geant4 in a set of simplified but increasingly challenging phantom geometries.
Approach: Building on previous work in Crossley et al. (2025), where energy deposition from a proton beam was modelled using an SDE framework, we implemented the model using standard approximations to interaction cross sections and mean excitation energies, enabling straightforward adaptation to new materials and configurations. The model was benchmarked against Geant4 in homogeneous, longitudinally heterogeneous and laterally heterogeneous phantoms, for assessment of depth-dose behaviour, lateral transport and impact of material heterogeneities.
Main results: Across all phantom configurations and beam energies, the SDE model reproduced the main depth–dose characteristics predicted by Geant4, with proton range agreement within 0.2 mm for 100 MeV beams and within 0.6 mm for 150 MeV beams. Voxel-wise comparisons yielded gamma pass rates exceeding 95% for all cases under strict 2%/0.5 mm criteria with a 1% dose threshold. Differences between the two approaches were spatially localised and primarily associated with regions of steep dose gradients or material heterogeneities, while overall lateral beam dispersion was well reproduced. In terms of computational performance, the SDE model achieved speed-up factors of approximately 2.5–3 relative to single-threaded Geant4, consistently across different Geant4 physics lists.
Significance: These results demonstrate that the SDE-based approach can reproduce key dosimetric features predicted by high-fidelity Monte Carlo simulations with good accuracy while already offering a moderate reduction in computational cost. Owing to its formulation, the method is naturally amenable to parallel and GPU-accelerated implementations, suggesting potential for substantial further speed improvements. This makes the approach a promising candidate for fast dose calculations in proton therapy.
@@ 1 Introduction
Accurate dose calculations are critical for proton therapy treatment planning to ensure precise treatment delivery while minimising uncertainties in patient outcomes. Thus, high-fidelity models are needed to capture the inherently stochastic nature of proton interactions in tissue, including energy loss, scattering, and nuclear interactions. Currently, pencil beam algorithms (Hong et al., 1996; Schaffner et al., 1999) remain widely used in clinical workflows due to their computational efficiency, particularly for treatment planning and optimisation. However, they achieve speed at the expense of accuracy. Pencil beam algorithms perform reliably in homogeneous media and provide good agreement for depth–dose profiles, but struggle to accurately model lateral spread, tissue heterogeneities, and complex geometries. Due to this, Monte Carlo (MC) models such as Geant4 (Agostinelli et al., 2003), TOPAS (Perl et al., 2012), GATE (Grevillot et al., 2020) and FLUKA (Böhlen et al., 2014), among others, are an alternative to analytical methods to perform the most accurate dose calculations, which can reduce the error margins by several millimetres (Paganetti, 2012). This is achieved by tracking each proton step by step and sampling each possible interaction according to its probability, which means this method tends to be much slower, compromising computational efficiency and therefore its suitability for routine treatment planning.
Optimised CPU codes such as MCsquare (Deng et al., 2020) leverage multi-core architectures to achieve substantial reductions in computation time. To date, a number of highly optimised GPU-based Monte Carlo frameworks have demonstrated that near real-time proton dose calculation is achievable in practice, including FRED (Schiavi et al., 2017), pGPUMCD (Maneval et al., 2019) and gPMC (Jia et al., 2012). Other GPU-based implementations such as MOQUI (Lee et al., 2022) enable highly efficient dose scoring, while other approaches avoid simulating secondary particles to fully exploit GPU computing power (Shan et al., 2022). These implementations exploit massive parallelism to dramatically reduce wall-clock time. These methods are increasingly being adopted in research and clinical workflows (Feng et al., 2022).
The present work does not aim to compete with or replace these GPU-based Monte Carlo solutions in terms of absolute runtime. Instead, the focus here is on demonstrating that, proton-for-proton and for a given level of dosimetric accuracy, the proposed SDE formulation can reduce the computational effort required per proton to obtain a dose estimate compared to conventional particle-tracking Monte Carlo. In this sense, the SDE framework should be viewed as a complementary modelling paradigm: it preserves the stochastic physics of transport while replacing frequent small-scale interactions with an analytically motivated diffusion process, thereby improving efficiency at the level of the underlying physical model rather than through hardware acceleration alone. This provides a mathematically distinct and potentially synergistic alternative that, due to the inherent independence of proton paths in the SDE model, can be integrated with modern computing architectures in future implementations.
Rather than relying exclusively on increasing computational power, recent efforts have also focused on advancing the underlying physical and mathematical models to improve efficiency at the algorithmic level. In addition to variance reduction techniques such as the Leq formalism and related approaches (1), which accelerate Monte Carlo simulations by reducing statistical noise, there is growing interest in alternative formulations of the transport problem itself.
@@
In this context, new mathematical approaches are needed for modelling proton beam treatment delivery, aiming to lie between high-fidelity Monte Carlo accuracy and pencil beam algorithm efficiency. Ashby et al. (2024) proposed an approach based on optimisation by computing the fluence of the associated Boltzmann transport equation numerically. In Crossley et al. (2025), a complementary approach was introduced to solve the Boltzmann transport equation in backward form using a jump stochastic differential equation (SDE). The latter retains the probabilistic nature of Monte Carlo simulations while replacing frequent small-angle Coulomb scatterings with a physics-informed diffusion term. This reduces computational cost without sacrificing physical fidelity due to the central limit theorem. Moreover, Kyprianou et al. (2025) showed that the two methods are consistent in the sense that they are alternative but compatible mathematical descriptions of the same underlying dynamics, expressed in terms of macroscopic fluence on the one hand and microscopic particles on the other. The reduced runtime of the proposed SDE approach compared to full Monte Carlo simulations makes it an attractive candidate for tasks such as plan verification, robustness evaluation, and adaptive planning, where high accuracy is required but computational cost remains a limiting factor.
@@
  • To show SDE-based methods provide accurate dose predictions at low computational cost per proton in clinically relevant settings.
@@
2 Methods
2.1 SDE-based transport formulation
High-energy protons decelerating via interactions with subatomic particles are the basis of proton beam therapy. Deceleration is directly associated with energy deposition, and this energy transfer is a key component of attacking cancerous tissues. The rate of interaction of a proton with these subatomic particles increases as it loses energy, which leads to greater energy deposition towards the end of a proton track. Hence, the position at which a proton beam deposits the majority of its energy is controllable, limiting the exposure of surrounding healthy tissues. Indeed, the energy deposition can be arranged in such a way that there is effectively no exit dose from a proton beam directed towards tumour tissues. Proton beam therapy is distinguished from photon therapy by these phenomena. In Crossley et al. (2025), the basic principles of nuclear physics were used to model the dynamics of protons travelling through matter in three-dimensional space. The principal mathematical tool was a so-called stochastic differential equation (SDE) with jumps. The SDE method lends itself well to new Monte Carlo approaches which favourably balance accuracy and computational cost. In this section, we begin with a brief reminder of proton beam physics and how it plays into the modelling choices that lead to the SDE formulation, which is then described in detail. For a more detailed read on the background physics of proton beams, the reader is referred to Newhauser and Zhang (2015); Paganetti (2020) and Gottschalk (2018), among many possible sources.
2.1.1 Physical basis and scope
@@
We introduce the space-direction-energy phase space as a state space on which to describe proton beam dynamics. Let $D\subset\mathbb{R}^{3}$ be a closed, bounded and convex spatial domain, $\mathbb{S}_{2}$ be the unit sphere in $\mathbb{R}^{3}$, and $\mathcal{E}=[\mathtt{e}_{\mathtt{min}},\mathtt{e}_{\mathtt{max}}]\subset[0,\infty)$ be the interval of energies a proton can take. We define the energy-position-direction phase space as $\mathcal{C}=\mathcal{E}\times D\times\mathbb{S}_{2}.$
@@
To describe the dynamics of a proton at configuration $x=({\mathrm{\epsilon}},r,\omega)\in\mathcal{C}$, we consider transport as well as three classes of interactions with surrounding matter (See Figure 1). For further details, we refer the interested reader to Newhauser and Zhang (2015). The three classes are:
@@
  • Transport. The proton moves in a straight line in the direction $\omega$ in the absence of other interactions.
@@
  • Inelastic Coulomb interaction. We assume that a proton loses energy continuously via frequent collisions with orbital electrons. For an instantaneous configuration $x=({\mathrm{\epsilon}},r,\omega)$, there are two contributing factors to this continuous energy loss. Deterministic loss occurs at a rate given by the stopping cross section, $\varsigma_{1}(x)$, per unit track length. We model variability in the number of orbital electron collisions per unit track length by an additional stochastic loss term, governed by a Brownian motion whose volatility is given by the energy-straggling cross section $\varsigma_{2}(x)$. We assume that these collisions have no effect on the direction of motion of the relatively massive proton, or equivalently, that any effect is subsumed into the elastic scatter term described below.
@@
    Figure 1: The three main interactions of a proton with its surrounding matter. An elastic scattering (top) with a nucleus, a proton-nucleus collision which may be elastic or inelastic (centre), and an inelastic Coulomb interaction with an orbital electron (bottom).
@@
  • Elastic scatter. This class of events corresponds to changes in a proton’s direction due to an interaction with a nearby nucleus whilst conserving energy between the proton and the nucleus. The three types of elastic scattering events we consider are as follows:
@@
    Elastic Coulomb scatter: The proton passes close enough to a nucleus to feel a repulsive positive charge, resulting in a change in its direction of motion. We distinguish between large and small scatters as separate cases. For incoming direction $\omega$ and outgoing direction $\omega^{\prime}$, let $\theta\in[0,\pi]$ denote the polar scattering angle. We fix a small cut-off $\delta>0$, such that $\theta>\delta$ is a large (elastic Coulomb) scatter while $\theta\leq\delta$ is a small scatter. A cross section $\sigma_{\rm e}(x)\pi_{{\rm e}}(x;{\mathrm{d}}\omega^{\prime})$ specifies large scatter events, where $\sigma_{\rm e}(x)$ is the rate at which a proton $x=(\epsilon,r,\omega)\in\mathcal{C}$ undergoes a large elastic Coulomb scatter, while the density $\pi_{{\rm e}}(x;{\mathrm{d}}\omega^{\prime})$ describes the distribution of the resulting outgoing direction $\omega^{\prime}$. Small scatter events are subsumed into an aggregated term, described below. It is common to think of $\sigma_{\rm e}(x)$ as a rate per unit track length, capturing the number nuclear interactions a proton has with its surrounding medium.
@@
    Elastic proton-nucleus collision and scatter: In rare cases, the incoming proton is absorbed into a nucleus, which becomes excited and releases another proton with the same energy as the incoming one. Since protons are indistinguishable, we model them identically to the elastic Coulomb scatters described above, so that $\sigma_{e}$ and $\pi_{e}$ are the cross section and probability density of both event types combined.
@@
    Coulomb-nuclear interference: It is insufficient to treat the effects of elastic Coulomb scattering and elastic proton-nucleus collisions independently (Nekrasov, 2024). The combination of these effects results in an additional perturbation in the outgoing angle due to so-called Coulomb-nuclear interference. We take the same cut-off $\delta$ to distinguish between small and large scatters, where large scatters are subsumed into the elastic Coulomb scatter and small scatters subsumed into the aggregation of small scatter, described below. Mathematically, $\sigma_{e}$ and $\pi_{e}$ incorporate Coulomb-nuclear interference as well.
@@
  • Aggregated small scatter. We model the net effect of all small scatter events via a diffusive component in the direction of the motion $\omega$ of a proton $x=(\epsilon,r,\omega)\in\mathcal{C}$ (Vassiliev, 2017, Section 3.9.6). More precisely, the direction of transport $\omega$ undergoes a Brownian motion on $\mathbb{S}_{2}$ with state-dependent volatility $m(x)\geq 0$. We assume no energy loss is associated with these small scattering events.
@@
  • Inelastic proton-nucleus collision and scatter. When a proton interacts with a nucleus, it can undergo an inelastic collision, forming an excited nuclear state. This excited nucleus may subsequently emit a spray of secondary particles, transferring some of its energy to each particle. For the SDE model considered in this article, we assume that each collision will emit exactly one secondary proton which carries the majority of the outgoing energy and has outgoing direction $\omega^{\prime}$. The energy deficit between the incoming and outgoing protons is accounted for by the recoil of the nucleus, other secondary particle emission which our SDE neglects, and the binding energy of the interaction. Nuclear inelastic scattering occurs with cross section, or rate per unit track length $\sigma_{{\rm ne}}(x)$. At an inelastic scattering event, the configuration $x=(\epsilon,r,\omega)$ of the incoming proton is transformed to the outgoing configuration $x^{\prime}=({\mathrm{\epsilon}}(1-u),r,\omega^{\prime})$ with probability density $\pi_{{\rm ne}}(x;{\mathrm{d}}\omega^{\prime},{\mathrm{d}}u)$, where $u\in(0,1]$.
@@
Together, these interactions manifest in what is known as the Bragg peak: a graphical representation of the energy deposition per unit length along the axial direction of a proton beam. A typical Bragg peak curve is proportional to the average energy deposition per unit length along sequential proton tracks. Its characteristic shape shows a gradual increase in energy deposition followed by a sharp rise near the end of the proton range, and then a rapid fall-off. One of the defining features of the Bragg peak is its range, defined as the depth at which the dose falls to a predefined fraction of its maximum. Quantities such as R90 and R50, defined as the depths at which the dose falls to 90% and 50% of the maximum, respectively, are commonly used in the literature. In the context of the present work, the notion of a sequential proton track was introduced in Crossley et al. (2025) to describe the configuration-space dynamics of a proton, concatenated with subsequent protons that continue its trajectory after proton-nuclear interactions, ultimately to absorption or a de-energised state. Each sequential proton track is a random trajectory and a proton beam is comprised of many such tracks, all superimposed. We assume all sequential proton tracks are independent and hence, by the Law of Large Numbers, a proton beam is the average effect a sequential proton track multiplied by the number of protons in the beam.
2.1.2 Sequential proton track SDE formulation